# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v2

@@ 3.1 Condition for EWN
where the right hand side defines the additional bounds $\Delta t_{i,EWN}$ [^11] Among the two bounds the $i=1$ bound is more stringent. However, as we will soon show, satisfying $i=1$ bound is not necessary to satisfy EWN. Although the $i=1$ constraint is not relevant when formulating necessary and sufficient conditions for EWN it will reappear in our discussion of trying to understand non-locality in cutoff holography in Section 3.4..
@@
Quick inspection of the expression in Eq. (42) reveals that $\Delta t_{c}|_{\theta_{0}=\pi/2}=\Delta t_{A,EWN}|_{\theta_{0}=\pi/2}$. In other words, in the critical limit we recover standard holography results that EWN trivially follows from requiring spacelike separation of the intervals $A$ and $B$.
@@ 3.3 EWN and Spacelike Separation of Extremal Surfaces
Now we turn to the issue of understanding when $\chi_{1}$ and $\chi_{2}$ are spacelike separated. We start by claiming that anytime $|\Delta t|\geq\Delta t_{2,EWN}$ then $\chi_{1}$ and $\chi_{2}$ cannot be spacelike separated. To prove this, it is useful to make use of the fact that we can can express $\chi_{2}$ as the intersection of lightcones located at the boundary whose apexes are located at the points given in Eq. (27). [^15] This statement is strictly speaking only true when $|k_{2}|\leq 1$. However, note that if $|\Delta t|=\Delta t_{2,EWN}$ we have that $k_{2,EWN}=\frac{\Delta t_{2,EWN}}{a_{2}+b_{2}\sin\theta_{0}}=1+\frac{b_{1}+b_{2}\cos(2\theta_{0})-\sqrt{(b_{1}+b_{2}\cos(2\theta_{0}))^{2}+b_{2}^{2}\sin^{2}(2\theta_{0})}}{2\sin\theta_{0}(a_{2}+b_{2}\sin\theta_{0})}$¡1 when $\theta_{0}\in(0,\pi/2)$. For the purposes of our proof we need only show $\chi_{1}$ and $\chi_{2}$ fail to be spacelike separated from each other at $|\Delta t|=\Delta t_{2,EWN}$ and also very slightly above the bound as well so there is no issue here. Furthermore, the numerical calculations we did in Appendix B.4 make no such assumption and agree with what we have proved using this method. This means that $\chi_{1}$ is spacelike separated to $\bar{\chi}_{2}$ (i.e. the extension of $\chi_{2}$ behind brane which we discussed near the end of Section 2.3) whenever [^16] Note the orientation of the inequality is correct. Since $\chi_{1}$ should be contained in the region enclosed by the lightcones whose apexes are at $(t_{\pm,2},x_{\pm,2},0)$ which makes $\chi_{1}$ timelike to $(t_{\pm,2},x_{\pm,2},0)$.
@@ 3.4 Consequences for Locality in Cutoff Holography
In standard holography the boundary theory is local and two subregions are independent if there are not causally related, i.e., they are spacelike separated. However, in cutoff holography one generally expects the theory on the cutoff to be a non-local theory and it is less clear how independent subregions can be defined. It has been known in our setup involving ETW branes that spacelike separation in the bulk sense is stronger than spacelike separation in the boundary sense Omiya:2021olc . An interesting open question in such setups is which notion of “spacelike” separation we need to define subregions on the holographic cutoff surface that are independent of each other. A naive candidate condition would be that the boundary subregions should be spacelike separated through the bulk as this ensures no causal signaling can occur between $A$ and $B$ either through the bulk or boundary.
@@
Our computation in Eq. (42) proves that this requirement is not sufficient. In particular, as we have seen, it is possible to construct configurations in which $A$ and $B$ cannot influence each other causally through the bulk, but the entanglement wedges can. Such examples cannot be constructed in the case where $\theta_{0}=\pi/2$ (i.e. in standard holography). This implies that — provided entanglement wedges work in cutoff holography in the same way they do in standard AdS/CFT — $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$, and thus $A$ and $B$, are not independent of each other even when $A$ and $B$ are spacelike separated through the bulk. We suggest that this should be regarded a geometric bulk manifestation of the non-local nature of the theory on the cutoff. In such cases our bound on entanglement wedge nesting disallows one to consider $A$ and $B$ as independent subregions. And clearly, the condition that $\mathcal{W}_{E}(A)$ cannot causally influence $\mathcal{W}_{E}(B)$ is a necessary condition for “independent” subregions in non-local holographic theories such as the one we have on the brane.
@@
In addition, it is also interesting to consider what the condition $|\Delta t|<\Delta t_{A,EWN}$ implies about the separation between $A$ and $\text{Vir}(B)$ (which was used in the discussion of Eq. (23)). Taking inspiration from recent works which think about a notion of induced causality on the brane in terms of subregions that exist behind the brane on the imaginary conformal boundary Mori:2023swn ; Franken:2024wmh , we will think of $B$ as a coarse grained version of $\text{Vir}(B)$ which lives on the imaginary conformal boundary behind the brane. Through this identification we might suggest a natural way to think about independent subregions on the brane/cutoff is to restrict $B$ to be the such that $\text{Vir}(B)$ is spacelike separated from $A$ in the full spacetime without the brane. Interestingly enough, this exactly gives the condition that $|\Delta t|<\Delta t_{1,EWN}$ which was derived as a sufficient condition for EWN in Eq. (36) [^19] Recall that in Section 3.1, this condition represented the most stringent sufficient condition to ensure $\mathcal{W}_{E}(B)\subseteq\mathcal{W}_{E}(A\cup B)$, we disregarded it because it was actually not a necessary condition but it reappears here as well.. Using the fact that $\Delta t_{2,EWN}-\Delta t_{1,EWN}=2a_{-}$ along with the identity in Eq. (47) we can easily see that that $\Delta t_{1,EWN}-\Delta t_{A,EWN}=-\ell_{*}\leq 0$, where saturation occurs when $\theta_{0}=\pi/2$. This implies that any time $A$ and $\text{Vir}(B)$ are spacelike separated then EWN is will be satisfied. However, due to the fact that $\Delta t_{1,EWN}-\Delta t_{A,EWN}\leq 0$ there also exist configurations where EWN is satisfied but $A$ and $\text{Vir}(B)$ fail to be spacelike separated and therefore would not be independent in this alternate notion of subregion independence formulated in terms of subregions behind the brane.
@@ 3.5 Relation to Restricted Maximin
Trying to define holography on a cutoff surface does not only cause problem for EWN, but also other inequalities which need to be obeyed by entanglement entropies, such as strong subadditivity (SSA). Based on ideas put forward in Wall:2012uf , the authors of Grado-White:2020wlb proposed a prescription, called restricted maximin, which associates a quantity which obeys SSA to any set of achronal co-dimension two bulk regions. This quantity is thus a natural candidate for a holographic entanglement entropy. It was further shown in Grado-White:2020wlb that their version of entanglement entropy also obeys monogamy of mutual information and the corresponding entanglement wedges satisfy EWN. It is interesting to contrast our construction of entanglement wedges to the construction using restricted maximin Grado-White:2020wlb .
@@
Figure 11: We fixed $a_{1}=1,a_{2}=5,b_{1}=7,b_{2}=10$ and vary $\theta_{0}$ between 00 and $\pi/2$. Our construction of all entanglement wedges is well defined below . Restricted maximin can be applied in the parameter space below the solid green line. Restricted maximin and our prescription are expected to give same RT surfaces in shaded blue region where EWN is respected.
@@ 4 Entanglement Wedge Nesting in a Planar BTZ Black Hole Geometry
In this section, we will study conditions for entanglement wedge nesting for a two-sided planar BTZ black hole geometry with one exterior having an ETW brane by adapting the formalisms developed in the previous sections in AdS3. BTZ black hole geometry has also been of great interest in a variety of contexts ranging from black hole microstates Hartman:2013qma ; Kourkoulou:2017zaj ; Miyaji:2021ktr , brane world cosmologies Cooper:2018cmb ; Antonini:2019qkt ; Waddell:2022fbn , as well as constraining DGP couplings in brane-world setups Lee:2022efh ; Geng:2022slq ; Geng:2022tfc ; Geng:2023qwm , which motivates their study using our techniques.
In Section 4.1 we will introduce the two-sided BTZ black hole along with the ETW brane which will cut off a portion of the right exterior. We will also describe the configuration of intervals we will consider on the two exteriors. In particular, $A_{\text{Bdry}}$ is an interval defined on the conformal boundary in the left exterior and $A_{\text{Br}}$ is an interval defined on the right exterior on the brane. After doing this in Section 4.2 we derive expressions for the extremal surfaces we will be considering, which will be anchored to $A_{\text{Bdry}}$ and $A_{\text{Br}}$. In 4.2.1 we give explicit expressions for thermal RT surfaces anchored to $A_{\text{Bdry}}$ and $A_{\text{Br}}$ which remain in their respective exteriors. In 4.2.2 we compute the connected RT surface which goes through the horizon and connects $A_{\text{Bdry}}$ and $A_{\text{Br}}$. In Section 4.3 we give explicit expressions characterizing the wedges associated with the thermal RT surfaces depicted in Figures 13(a) - 13(b). In Section 4.4 we show that EWN of $\mathcal{W}_{E}(A_{\text{Br}})$ constrains the relative time shift between $A_{\text{Bdry}}$ and $A_{\text{Br}}$ in a way that prevents the RT surface in the connected phase from crossing the black hole “singularity”.
@@ 4.1 Two-Sided BTZ Black Holes with an ETW Brane
We start by discussing two-sided planar BTZ black hole with one side having an ETW brane. Nothing in what follows will depend on the fact that the black hole is planar, as long as the horizon size is sufficiently big relative to the brane and boundary intervals. The line element for the two sided BTZ black hole written in Kruskal-Szekeres coordinates is
@@ 4.2 Subregions and RT Surfaces
For the sake of mathematical simplicity we will restrict ourselves to a very limited configuration of boundary and brane intervals. In particular, we will consider constant Schwarzschild time slice intervals of equal length on both the boundary and brane centered at $x=0$. We will only allow for a relative shift in time for the boundary and brane intervals. We will denote the the constant time interval on the left boundary at Schwarzschild time $t=t_{\text{Bdry}}$ as $A_{\text{Bdry}}$. The constant time interval on the brane in the right exterior at Schwarzschild time $t=t_{\text{Br}}$ will be denoted by $A_{\text{Br}}$. For the configuration of intervals on the asymptotic boundary and brane described above there are two kinds of possible RT surfaces which are discussed in the next two subsections.
@@ 4.2.1 Thermal Extremal Surfaces
Let us begin by finding extremal surfaces which stay within a single exterior region; we will call those thermal extremal surfaces. In the BTZ Due to time-translation symmetry in each region, thermal extremal surfaces are located on a constant Schwarzschild time-slice. We denote the thermal surface in left exterior anchored to the boundary as $\chi_{\text{ther}}(A_{\text{Bdry}})$ and the thermal surface in the right exterior anchored to the brane as $\chi_{\text{ther}}(A_{\text{Br}})$. The extremal surfaces $\chi_{\text{ther}}(A_{\text{Bdry}})$ and $\chi_{\text{ther}}(A_{\text{Br}})$ are analogous to $\chi_{\text{dis}}(A)$ and $\chi_{\text{dis}}(B)$ in our previous AdS3 setup which will be used to construct disconnected wedges in our study of EWN. Thermal extremal surfaces extremize the area functional which in the BTZ geometry reads
@@
Using the expression above for the thermal RT surface anchored to a boundary interval we can also easily deduce the expression for the thermal RT surface which is anchored to a constant time slice interval on the brane at time $t=t_{\text{Br}}$ in the right exterior (i.e. $\chi_{\text{ther}}(A_{\text{Br}})$). To do this we define the “virtual interval” associated to the interval on the brane denoted $\text{Vir}(A_{\text{Br}})$ which is defined to be an interval which is obtained by continuing the RT surface anchored to the brane past the brane into the non-physical region. We can characterize this interval below,
@@
The thermal RT surface $\chi_{\text{ther}}(\text{Vir}(A_{\text{Br}}))$ is then described by
@@
We can use our results to rewrite the same thermal RT surface in Kruskal coordinates using the inverse of the transformation Eq. (58). With a little bit of algebra and a careful distinction between the left and right exteriors the transformation from Schwarzschild to Kruskal coordinates reads

### v3

@@ 3.1 Condition for EWN
where the right hand side defines the additional bounds $\Delta t_{i,EWN}$. [^12] Among the two bounds the $i=1$ bound is more stringent. However, as we will soon show, satisfying $i=1$ bound is not necessary to satisfy EWN. Although the $i=1$ constraint is not relevant when formulating necessary and sufficient conditions for EWN it will reappear in our discussion of trying to understand non-locality in cutoff holography in Section 3.4.
@@
Quick inspection of the expression in Eq. (42) reveals that $\Delta t_{c}|_{\theta_{0}=\pi/2}=\Delta t_{A,EWN}|_{\theta_{0}=\pi/2}$. In other words, in the critical limit we recover standard holography results that EWN trivially follows from requiring spacelike separation of the intervals $A$ and $B$ [^14] One can also take the parameters used to generate Figures 7,8,9 and verify that those choices of parameters do indeed satisfy Eq. (43) which is consistent with the features we described in those example plots.
@@ 3.3 EWN and Spacelike Separation of Extremal Surfaces
Now we turn to the issue of understanding when $\chi_{1}$ and $\chi_{2}$ are spacelike separated. We start by claiming that anytime $|\Delta t|\geq\Delta t_{2,EWN}$ then $\chi_{1}$ and $\chi_{2}$ cannot be spacelike separated. To prove this, it is useful to make use of the fact that we can express $\chi_{2}$ as the intersection of lightcones located at the boundary whose apexes are located at the points given in Eq. (27). [^17] This statement is strictly speaking only true when $|k_{2}|\leq 1$. However, note that if $|\Delta t|=\Delta t_{2,EWN}$ we have that $k_{2,EWN}=\frac{\Delta t_{2,EWN}}{a_{2}+b_{2}\sin\theta_{0}}=1+\frac{b_{1}+b_{2}\cos(2\theta_{0})-\sqrt{(b_{1}+b_{2}\cos(2\theta_{0}))^{2}+b_{2}^{2}\sin^{2}(2\theta_{0})}}{2\sin\theta_{0}(a_{2}+b_{2}\sin\theta_{0})}$¡1 when $\theta_{0}\in(0,\pi/2)$. For the purposes of our proof we need only show $\chi_{1}$ and $\chi_{2}$ fail to be spacelike separated from each other at $|\Delta t|=\Delta t_{2,EWN}$ and also very slightly above the bound as well so there is no issue here. Furthermore, the numerical calculations we did in Appendix B.4 make no such assumption and agree with what we have proved using this method. This means that $\chi_{1}$ is spacelike separated to $\bar{\chi}_{2}$ (i.e. the extension of $\chi_{2}$ behind brane which we discussed near the end of Section 2.3) whenever [^18] Note the orientation of the inequality is correct. Since $\chi_{1}$ should be contained in the region enclosed by the lightcones whose apexes are at $(t_{\pm,2},x_{\pm,2},0)$ which makes $\chi_{1}$ timelike to $(t_{\pm,2},x_{\pm,2},0)$.
@@ 3.4 Consequences for Locality in Cutoff Holography
In standard holography the boundary theory is local and two subregions are independent if there are not causally related, i.e., they are spacelike separated. However, in cutoff holography one generally expects the theory on the cutoff to be a non-local theory and it is less clear how independent subregions can be defined. It has been known in our setup involving ETW branes that spacelike separation in the bulk sense is stronger than spacelike separation in the boundary sense due to shortcuts that exist through the bulk between points on the brane and boundary Omiya:2021olc . [^21] As a note to the reader it was suggested in Omiya:2021olc that the existence of such shortcuts was in tension with having a causal effective field theory (EFT) in the brane-boundary system (i.e. the “intermediate description”). In the recent work Geng:2025yys , it was shown that this is not problematic when one views the theory in the intermediate description as an EFT with a cutoff. It was argued that the geodesic shortcut is actually a UV effect from the perspective of the brane-boundary system (it is encoded by heavy modes KK modes). Since the EFT on the brane-boundary system comes with a cutoff it is safe from this UV effect. Coupling to the heavy KK modes in the UV was also explored as they could potentially lead to signatures of causality violation, but in this case it was shown that constraints from unitarity also prevented non-causal signatures from appearing in correlators. An interesting open question in such setups is which notion of “spacelike” separation we need to define subregions on the holographic cutoff surface that are independent of each other. A naive candidate condition would be that the boundary subregions should be spacelike separated through the bulk as this ensures no causal signaling can occur between $A$ and $B$ either through the bulk or boundary.
@@
Our computation in Eq. (42) proves that this requirement is not sufficient. In particular, as we have seen, it is possible to construct configurations in which $A$ and $B$ cannot influence each other causally through the bulk, but the entanglement wedges can. Such examples cannot be constructed in the case where $\theta_{0}=\pi/2$ (i.e. in standard holography). This implies that — provided entanglement wedges work in cutoff holography in the same way they do in standard AdS/CFT — $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$, and thus $A$ and $B$, are not independent of each other even when $A$ and $B$ are spacelike separated through the bulk. We suggest that this might be regarded a geometric bulk manifestation of the non-local nature of the theory on the cutoff. In such cases our bound on entanglement wedge nesting disallows one to consider $A$ and $B$ as independent subregions. And clearly, the condition that $\mathcal{W}_{E}(A)$ cannot causally influence $\mathcal{W}_{E}(B)$ is a necessary condition for “independent” subregions in non-local holographic theories such as the one we have on the brane. Here we are not making any claims in contradiction/tension to the findings of Geng:2025yys . The existence of such non-local correlations between $A$ and $B$ do not pose a problem to the causality of the EFT on the brane-boundary system. Since the effect described here is geometric and in the bulk, we expect these to be UV effects which manifest well above the cutoff of the EFT, thereby leaving the local EFT structure at low energies intact. Similar to how the geodesic shortcut in bulk was argued in Geng:2025yys to be encoded by heavy KK modes in the UV regime we might speculate a similar description also exists for the enhanced non-local correlations between $A$ and $B$ described here. We leave more detailed investigations in such a direction to future work.
@@
To conclude the discussion of this section it is also interesting to consider what the condition $|\Delta t|<\Delta t_{A,EWN}$ implies about the separation between $A$ and $\text{Vir}(B)$ (which was used in the discussion of Eq. (23)). Taking inspiration from recent works which think about a notion of induced causality on the brane in terms of subregions that exist behind the brane on the imaginary conformal boundary Mori:2023swn ; Franken:2024wmh , we will think of $B$ as a coarse grained version of $\text{Vir}(B)$ which lives on the imaginary conformal boundary behind the brane. Through this identification we might suggest a natural way to think about independent subregions on the brane/cutoff is to restrict $B$ to be the such that $\text{Vir}(B)$ is spacelike separated from $A$ in the full spacetime without the brane. Interestingly enough, this exactly gives the condition that $|\Delta t|<\Delta t_{1,EWN}$ which was derived as a sufficient condition for EWN in Eq. (36). [^22] Recall that in Section 3.1, this condition represented the most stringent sufficient condition to ensure $\mathcal{W}_{E}(B)\subseteq\mathcal{W}_{E}(A\cup B)$, we disregarded it because it was actually not a necessary condition but it reappears here as well. Using the fact that $\Delta t_{2,EWN}-\Delta t_{1,EWN}=2a_{-}$ along with the identity in Eq. (47) we can easily see that $\Delta t_{1,EWN}-\Delta t_{A,EWN}=-\ell_{*}\leq 0$, where saturation occurs when $\theta_{0}=\pi/2$. This implies that any time $A$ and $\text{Vir}(B)$ are spacelike separated then EWN is will be satisfied. However, due to the fact that $\Delta t_{1,EWN}-\Delta t_{A,EWN}\leq 0$ there also exist configurations where EWN is satisfied but $A$ and $\text{Vir}(B)$ fail to be spacelike separated and therefore would not be independent in this alternate notion of subregion independence formulated in terms of subregions behind the brane.
@@ 3.5 Relation to Restricted Maximin
Trying to define holography on a cutoff surface does not only cause problems for EWN, but also other inequalities which need to be obeyed by entanglement entropies, such as strong subadditivity (SSA). Based on ideas put forward in Wall:2012uf , the authors of Grado-White:2020wlb proposed a prescription, called restricted maximin, which associates a quantity which obeys SSA to any set of achronal co-dimension two bulk regions. This quantity is thus a natural candidate for a holographic entanglement entropy. It was further shown in Grado-White:2020wlb that their version of entanglement entropy also obeys monogamy of mutual information and the corresponding entanglement wedges satisfy EWN. It is interesting to contrast our construction of entanglement wedges to the construction using restricted maximin Grado-White:2020wlb .
@@
Figure 11: We fixed $a_{1}=1,a_{2}=5,b_{1}=7,b_{2}=10$ and vary $\theta_{0}$ between 00 and $\pi/2$. Our construction of all entanglement wedges is well defined below $\min\{\Delta t_{2,EWN},\Delta t_{c}\}$ which is the union of the gray and blue shaded regions. Restricted maximin can be applied in the parameter space below the solid green line. Restricted maximin and our prescription are expected to give same RT surfaces in shaded blue region where EWN is respected.
@@ 4 Entanglement Wedge Nesting in a Planar BTZ Black Hole Geometry
In this section, we will study conditions for entanglement wedge nesting for a two-sided planar BTZ black hole geometry with one exterior having an ETW brane. We will adapt the formalisms developed for AdS3 in the previous sections and apply them to the BTZ black hole. The BTZ black hole geometry has been of great interest in a variety of contexts ranging from black hole microstates Hartman:2013qma ; Kourkoulou:2017zaj ; Miyaji:2021ktr , brane world cosmologies Cooper:2018cmb ; Antonini:2019qkt ; Waddell:2022fbn , as well as constraining DGP couplings in brane-world setups Lee:2022efh ; Geng:2022slq ; Geng:2022tfc ; Geng:2023qwm , which motivates their study using our techniques.
In Section 4.1 we will introduce the two-sided BTZ black hole along with the ETW brane which will cut off a portion of the right exterior. We will also describe the configuration of intervals we will consider in the two exteriors. In particular, $A_{\text{Bdry}}$ is an interval defined on the conformal boundary in the left exterior and $A_{\text{Br}}$ is an interval defined in the right exterior on the brane. After doing this, in Section 4.2 we derive expressions for the extremal surfaces we will be considering which will be anchored to $A_{\text{Bdry}}$ and $A_{\text{Br}}$. In 4.2.1 we give explicit expressions for thermal RT surfaces anchored to $A_{\text{Bdry}}$ and $A_{\text{Br}}$ which remain in their respective exteriors. In 4.2.2 we compute the connected RT surface which goes through the horizon and connects $A_{\text{Bdry}}$ and $A_{\text{Br}}$. In Section 4.3 we give explicit expressions characterizing the wedges associated with the thermal RT surfaces depicted in Figures 13(a) - 13(b). In Section 4.4 we show that EWN of $\mathcal{W}_{E}(A_{\text{Br}})$ constrains the relative time shift between $A_{\text{Bdry}}$ and $A_{\text{Br}}$ in a way that prevents the RT surface in the connected phase from crossing the black hole “singularity”.
@@ 4.1 Two-Sided BTZ Black Holes with an ETW Brane
We start by discussing the two-sided planar BTZ black hole with one side having an ETW brane. Nothing in what follows will depend on the fact that the black hole is planar, as long as the horizon size is sufficiently big relative to the brane and boundary intervals. The line element for the two sided BTZ black hole written in Kruskal-Szekeres coordinates is
@@ 4.2 Subregions and RT Surfaces
For the sake of mathematical simplicity we will restrict ourselves to a very limited configuration of boundary and brane intervals. In particular, we will consider constant Schwarzschild time slice intervals of equal length on both the boundary and brane centered at $x=0$. We will only allow for a relative shift in time for the boundary and brane intervals. We will denote the constant time interval on the left boundary at Schwarzschild time $t=t_{\text{Bdry}}$ as $A_{\text{Bdry}}$. The constant time interval on the brane in the right exterior at Schwarzschild time $t=t_{\text{Br}}$ will be denoted by $A_{\text{Br}}$. For the configuration of intervals on the asymptotic boundary and brane described above there are two kinds of possible RT surfaces which are discussed in the next two subsections.
@@ 4.2.1 Thermal Extremal Surfaces
Let us begin by finding extremal surfaces which stay within a single exterior region; we will call those thermal extremal surfaces. Due to time-translation symmetry in each region, thermal extremal surfaces are located on a constant Schwarzschild time-slice. We denote the thermal surface in left exterior anchored to the boundary as $\chi_{\text{ther}}(A_{\text{Bdry}})$ and the thermal surface in the right exterior anchored to the brane as $\chi_{\text{ther}}(A_{\text{Br}})$. The extremal surfaces $\chi_{\text{ther}}(A_{\text{Bdry}})$ and $\chi_{\text{ther}}(A_{\text{Br}})$ are analogues to $\chi_{\text{dis}}(A)$ and $\chi_{\text{dis}}(B)$ in our previous AdS3 setup which will be used to construct disconnected wedges in our study of EWN. Thermal extremal surfaces extremize the area functional which in the BTZ geometry reads
@@
Using the expression above for the thermal extremal surface anchored to a boundary interval we can also easily deduce the expression for the thermal extremal surface which is anchored to a constant time slice interval on the brane at time $t=t_{\text{Br}}$ in the right exterior (i.e. $\chi_{\text{ther}}(A_{\text{Br}})$). To do this we define the “virtual interval” associated to the interval on the brane denoted $\text{Vir}(A_{\text{Br}})$ which is defined to be an interval which is obtained by continuing the extremal surface anchored to the brane past the brane into the non-physical region. We can characterize this interval below,
@@
The thermal extremal surface $\chi_{\text{ther}}(\text{Vir}(A_{\text{Br}}))$ is then described by
@@
We can use our results to rewrite the same thermal extremal surface in Kruskal coordinates using the inverse of the transformation Eq. (58). With a little bit of algebra and a careful distinction between the left and right exteriors the transformation from Schwarzschild to Kruskal coordinates reads